OpenStudy (anonymous):

Write in order from least to greatest, 30%, 0.24, 2/5.

OpenStudy (anonymous):

We need to write each number either as a percentage, decimal or fraction. Which is easiest for you to do?

OpenStudy (anonymous):

i get it now thanks for ur help

OpenStudy (anonymous):

30/100 is 0.3 so it will be 0.24,30%,2/5

OpenStudy (ash2326):

sorry:( 30%=30/100=0.3 0.24= 0.24 2/5=0.4 from least to highest 0.24, 30%, 2/5.

OpenStudy (anonymous):

thanks